Solve the proportion please:
5/6 = a-2/4

Found 2 solutions by xcentaur, JimMarshall:
Answer by xcentaur(357)

5/6 = a-2/4


5(4) = 6(a-2)


20 = 6a - 12


20+12 = 6a


6a=32


a=32/6
a=16/3


Ans:a=16/3


Answer by JimMarshall(82)

==>5/6 = a-2/4
==>5(4) = 6(a-2)
==?20 = 6a - 12
==?20+12 = 6a
==?6a=32
==?a=32/6
==>a=16/3 Ans